OpenStudy (anonymous):

Graph with a line going through points 0,2.5 and 4, 2.5. Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1 @Nnesha @ttb123456789

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

slope is 0. so y=b -1=3*0+b b=-4 y=-4

OpenStudy (anonymous):

but uh let me redo

OpenStudy (anonymous):

thats not rt

OpenStudy (anonymous):

thats not a opption